The refractive index of the glass of a prism for red light is 1.514 and for blue light is 1.523. Calculate the difference in the velocities of the red and blue light in the prism if the velocity of light in vacuum is 3 x 10^(5) km/second.


1
Expert's answer
2021-03-12T09:24:31-0500

Difference between velocity can be given as

"\Delta v=c(\frac{1}{n_1}-\frac{1}{n_2})"


"\Delta v=(3\times10^5)(\frac{1}{1.514}-\frac{1}{1.523}) =1171km/s"